OpenStudy (anonymous):

Irma is walking along a circular path around a lake she walks 3 miles and the radius of the circle is 3.5. how ling is the central angle that intersects her path? 49 degrees 60 degrees 98degrees 147 degrees none of the above

OpenStudy (anonymous):

\[\frac{ x }{ 360 }=\frac{ 3 }{ 2\pi*3.5}\] Solve for x